Plumbing: The plumber wants to cut a 24-ft. length of pipe into three sections. The largest piece should be twice the length of the middle piece, and the middle piece should be twice the length of the smallest piece. How long should each each piece be?
Answer by askmemath(368)

Let the smallest pipe length = L
The length of middle pipe = "the middle piece should be twice the length of the smallest piece" = 2L
And Length of Largest Pipe = "The largest piece should be twice the length of the middle piece" = 2 x 2L = 4L
Now total length of 3pieces = L + 2L + 4L = 7L
But we know that total length = 24
Therefore 7L = 24
or L = 24/7 which is the length of the smallest pipe
Can you find out the length of the remaining pieces?